AliEn News

general:

25.09.07:

There will be a scheduled software upgrade on the CASTOR storage at 
CERN on 26/09/2007 from 08:00 hours. As a precaution, all user jobs will 
be suspended and relaunced after the end of the intervention. Some of the 
jobs may end in error due to proxy (TTL) expiration.
 
    Job submission in AliEn will not be suspended, however new jobs will 
not be scheduled to run before the end of the intervention.                

    The operation is scheduled to last ~8 hours

24.09.07:

  • Automatic histo merging in AliEn
Example 1: you want to merge your output histogram file 'histo.root' into 
a merge file 'histo-merged.root'. You just add this line to your master job:

Merge = { "histo.root:/alice/jdl/mergerootfile.jdl:histo-merged.root" };

=================================================================================
Example 2: you want to merge two different output files 'histo.root' & 
'results.root', then you just specify a list of merge jobs in the JDL:

Merge = { "histo.root:/alice/jdl/mergerootfile.jdl:histo-merged.root" , "results.root:/alice/jdl/mergerootfile.jdl:results-merged.root" };

21.09.06:

  • Alice::CERN::castor2 is not working
    • it is up again

26.09.06:

  • Connections to AliEn central services problematic. Network problems at CERN assumed

01.10.06

  • SRM endpoint "castorgrid.cern.ch" will be stopped.
    • replacement is srm.cern.ch
    • gridftp service on castorgrid.cern.ch will still be in use for some while

02.10.06 - general information

  • problems opening galice.root located on the Grid are related to
    • wrong behavior of TAlienFile. The problem is being fixed and will be committed to ROOT CVS soon

16.10.06

  • stoppage of voms.cern.ch as VOMS server
    • users have to be registered in lcg-voms.cern.ch
    • more info: see here

18.10.06

  • interruption of Castor service at CERN

20.10.06

  • GridKa will be shut down until at least Monday, 23.10.06

24.10.06

  • GridKa is operational again
  • CASTOR2 SE inaccessible from 2 pm to 3:30 pm today.

27.10.06

  • one of the xrootd data movers at CERN (disk server) is down for maintenance
  • dCache update to version 1.6.6-5 finished successfully at GridKa but gridftp- doors crash with unkown reason so far
  • The upgrade to dCache version 1.6.6-5 did not bring all the expected improvements. The stability of the gridftp doors has not improved.

30.10.06

  • xrootd server back into production

31.10.06

  • apiserver01, which is serving the user connections from aliensh, has stopped working (software) and cannot be restarted following the standard procedure. Still looking at the problem - more news later.

01.11.06

  • lcg-voms.cern.ch vomrs service will be unavailable for 15 mins at 10am UTC

03.11.06

  • CERN tape library will stop it''s activity TODAY at 14h30 and intervene for a period of ~1h30m.

08.11.06

  • A maintenance intervention on the CERN firewall will run on Wednesday the 8th of November 2006 between 6:00AM and 7:30AM CET (5:00-6:30 UTC).

15.11.06

  • A power dip at FZK at around 19:15 UTC has cut out some services.

24.11.06

  • PDC06 Grid production will come to a stop today. Central services stop at 08:30 am. FTS transfers will stop at 08:30 am, too. From 08:30 on no user will be able to login to AliEn until further notice. Regional services will be stopped, also.
  • ALICE wide upgrade to AliEn v2.12 will start in the order
    • central services
    • CERN job and storage services
    • sites
  • For more information on the update, please see the v.2-12 release notes at http://alien.cern.ch/twiki/bin/view/AliEn/Development

30.11.06

  • an intervention on one of the CERN Tape Libraries on Thursday this week (30/11/2006), starting at 09h00, is planned; the intervention should take 5h.
  • GGUS system will be upgraded to version 5.7.1. Therefore system will be down on 30th November from 14:00 to 16:00 UTC
  • From Thursday the 30th of November 7.30am to the 1st of December 7.30am, the CERN''s Internet traffic will transit over the new firewall system.
    • IMPACT: Some short instability in off-site network connectivity may be experienced during the switch to the new system, and when back to the current one.

04.12.06

  • On Monday morning there are CERN site wide power tests from 06.00 to 06.08. There should be no problems for computing services as the Computer centre will be protected by the UPS.
  • There will also be short cuts on Monday December 4th on the CERN Technical Network from 9.00 to 12.00; the General Network will not be affected.
  • The lcg-voms.cern.ch vomrs service will be unavailable for, we hope, a maximum period of 2 hours, on December 4th 2006 @16hrs CET
    • Reason: Upgrade to vomrs-1.3.0.

11.12.06

  • Due to the wide introduction of voms services, the LDAP based Authentication database will be stopped on Monday December 11th 2006 @10 a.m. hrs CET.

04.01.07

  • GGUS Downtime 2007-01-04 11:30-15:00 UTC

-- KilianSchwarz - 21 Sep 2006
Topic revision: r29 - 2007-09-25, KilianSchwarz
 
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding GSI Wiki? Send feedback
Imprint (in German)
Privacy Policy (in German)